# Constraining Participation: Affordances of Feedback Features in Interfaces to Large Language Models

### v1

@@ 6. Conclusion
Our analysis reveals that existing feedback features in the ChatGPT interface constrain participation around developers’ instrumental needs. This approach to user input and participation views engagement as a means to an end rather than an opportunity for dialogue. While deliberative approaches may address some limitations of existing mechanisms, the scope of user and non-user input remains constrained. To move beyond perfunctory engagement, we call for a shift in focus from user preference elicitation to building robust infrastructures for sustained, meaningful dialogue about the purpose and application of LLMs. As these technologies become increasingly embedded in information, workplace, and entertainment services, the urgency of this task cannot be overstated.
Appendix A Survey Results
Appendix A compiles the survey results from our study. It contains a series of tables with detailed breakdowns of respondent usage patterns, feedback behaviours, demographics, and perceptions of submitting feedback using the interactive features in the ChatGPT interface.
@@ Appendix A Survey Results
(Response rate: 526/526, Multiple selections allowed)
@@
About half the time I have used ChatGPT 98 24.5%
Most times I have used ChatGPT 92 23.0%
@@
Most times I have used ChatGPT 20 13.2%
About half the time I have used ChatGPT 17 11.2%
Every time I have used ChatGPT 3 2.0%
@@
Sometimes I have regenerated a response 120 36.3%
@@
Satisfied with the performance and have not felt the need to submit feedback 17 21.8%
@@
Contacting OpenAI (e.g., via email) 10 9.7%
@@
(Response rate 103/103, Multiple selections allowed)
@@
(Response rate: 509/526, up to 3 selections allowed)
Notes: ‘Other African languages’ includes Setswana, Xhosa, Yoruba, Swahili, and Hausa; ’Other European languages’ includes Finnish, Hungarian, Latvian, Russian, Norwegian, Danish, Swedish, Catalan, Ukrainian, and Lithuanian; ’Other Asian languages’ includes Hindi, Hebrew, Turkish, Arabic, Mongolian, Korean, Japanese, Persian, Bengali, Tagalog, Thai, and Sinhalese; ’Other languages’ includes Portuguese (Brazilian), Slovenian, Croatian, and unspecified others.
@@
(Response rate: 524/526)
Appendix B Statistical Relationships and Test Results
Appendix B presents our statistical analysis of relationships between key variables in our study. Table 31 presents the results of chi-square tests of associations between various factors such as user characteristics, feedback behaviours, and perceptions of submitting feedback through the ChatGPT interface.
Table 31. Chi-square Test Results
@@ Appendix B Statistical Relationships and Test Results
Thumbs-Up/down & Regenerating feedback 50.419 ¡.001 ¡.001* 1 .319 Mod. 497
Education/Training in AI & Confidence 12.773 .002 .023* 2 .160 Weak 497
Working in AI & Barriers 9.821 .002 .023* 1 .143 Weak 477
Region & Open-text feedback 13.990 .003 .023* 3 .192 Mod. 378
Prior AI use & Open-text feedback 8.926 .003 .023* 1 .154 Weak 378
Thumbs-Up/down & ChatGPT frequency 13.689 .008 .045* 4 .166 Weak 497
AI topics frequency & Confidence 20.726 .008 .045* 8 .144 Weak 497
Age & Confidence 17.068 .009 .045* 6 .131 Weak 497
Education/Training in AI & Open-text 6.132 .013 .059 1 .127 Weak 378
Gender & Regenerating feedback 10.427 .015 .061 3 .145 Weak 497
Age & Barriers 9.372 .025 .078 3 .140 Weak 477
Open-text & ChatGPT frequency 10.956 .027 .078 4 .170 Mod. 378
AI topics frequency & Open-text 10.937 .027 .078 4 .170 Mod. 378
Region & Regenerating feedback 9.151 .027 .078 3 .136 Weak 497
Working in AI & Open-text feedback 4.660 .031 .080 1 .111 Weak 378
Age & Regenerating feedback 8.678 .034 .080 3 .132 Weak 497
Prior AI use & Barriers 4.512 .034 .080 1 .097 Negl. 477
Education/Training in AI & Barriers 4.010 .045 .101 1 .092 Negl. 477
Gender & Confidence 12.513 .051 .103 6 .112 Weak 497
Prior AI use & Confidence 5.953 .051 .103 2 .109 Weak 497
Working in AI & Confidence 5.632 .060 .114 2 .106 Weak 497
Open-text & Regenerating feedback 3.424 .064 .117 1 .095 Negl. 378
AI topics frequency & Thumbs-Up/down 7.615 .107 .186 4 .124 Weak 497
Working in AI & Regenerating feedback 2.375 .123 .205 1 .069 Negl. 497
Regenerating & ChatGPT frequency 5.004 .287 .459 4 .100 Weak 497
Age & Open-text feedback 3.342 .342 .526 3 .094 Weak 378
Gender & Thumbs-Up/down 3.160 .368 .545 3 .080 Weak 497
AI topics frequency & Barriers 3.962 .411 .554 4 .091 Weak 477
AI topics frequency & Regenerating 3.932 .415 .554 4 .089 Weak 497
Prior AI use & Thumbs-Up/down 0.666 .415 .554 1 .037 Negl. 497
Gender & Barriers 2.482 .478 .607 3 .072 Weak 477
Education/Training in AI & Thumbs 0.468 .494 .607 1 .031 Negl. 497
Working in AI & Thumbs-Up/down 0.453 .501 .607 1 .030 Negl. 497
Gender & Open-text feedback 1.801 .615 .723 3 .069 Weak 378
Age & Thumbs-Up/down 1.713 .634 .725 3 .059 Negl. 497
Prior AI use & Regenerating feedback 0.140 .708 .773 1 .017 Negl. 497
Region & Barriers 1.361 .715 .773 3 .053 Negl. 477
Education/Training in AI & Regenerating 0.065 .799 .841 1 .011 Negl. 497
Region & Thumbs-Up/down 0.708 .871 .894 3 .038 Negl. 497
Region & Confidence 1.548 .956 .956 6 .039 Negl. 497
@@
  • $\chi^{2}$ test was not performed between Thumbs-Up/down & Open-text feedback. Open-text feedback was only available to users who provided Thumbs-Up/down feedback, making these variables dependent.
  • Table 32 provides a guide for interpreting Cramér’s V values, in accordance with (Cohen, 2013).

### v2

@@ 6. Conclusion
Our analysis of ChatGPT’s feedback features reveals significant constraints on user participation, based on a survey of early adopters’ experiences. The interface appears to encourage simple, frequent, and performance-focused feedback while discouraging collective input and discussions among users. We argue that this constrained approach to participation suggests limitations for meaningful dialogue about social and ethical issues regarding LLM deployment as system usage becomes more widespread.
Three research directions could substantiate and extend our analysis. Firstly, empirical studies could test whether deliberative processes better capture social and ethical concerns than current processes. Additionally, longitudinal research could track both how feedback patterns evolve as broader populations adopt LLMs, and how organisations frame their feedback processes over time. Such work might reveal growing divergences between companies’ stated commitments and actual practices. Most importantly, we encourage ethnographic and action research examining the practical work of infrastructuring. Ethnographic studies could document ongoing attempts to create participatory structures in specific domains, while action researchers could collaborate with affected communities to prototype and sustain new forms of engagement (see, e.g., (Tseng et al., 2025)).
@@
Without fundamental changes, the scope of user (and non-user) input will remain constrained. To move beyond perfunctory engagement requires shifting focus from user preference elicitation to building participatory infrastructures with affordances that allow and encourage sustained dialogue—including disagreement—with affected stakeholders about the purpose and applications of LLMs. As these technologies become increasingly embedded in information, workplace, and entertainment services, the urgency of this task cannot be overstated.
Acknowledgements.
This research is supported by an Australian Government Research Training Program (RTP) Scholarship and a Florence Violet McKenzie Scholarship. We thank the organisers and participants of the CHI 2023 Workshop on Designing Technology and Policy Simultaneously and the 2023 AusSTS Conference for feedback on earlier drafts. We also thank Professor Alice Richardson of the Statistical Support Network, ANU, for helpful advice on the statistical analysis in this article. Finally, we thank our reviewers for their feedback and suggestions.
Appendices
Appendix A Survey Questions
Appendix A lists the substantive questions from the online survey instrument used for this study.
Screening
  1. (1)
    Have you used ChatGPT in any capacity? e.g., to generate text, answer a question, or simply experiment with the system.
    • Yes
    • No
ChatGPT Usage
  1. (2)
    How often have you used ChatGPT since its release in November 2022?
    • Daily
    • A few times a week
    • Once a week
    • Monthly
    • Rarely
  2. (3)
    In which of the following ways have you used ChatGPT? Please select all that apply.
    • Academic research assistance (e.g., generating citations, summarising articles)
    • Generating text (e.g., drafting e-mails, writing essays or reports)
    • Answering specific questions (e.g., factual, technical, or theoretical questions)
    • Creative writing (e.g., stories, poetry, or brainstorming ideas)
    • Language learning or translation
    • Personal advice or emotional support (e.g., relationship or career advice)
    • Professional or technical problem-solving (e.g., programming, troubleshooting)
    • Entertainment or casual conversation (e.g., jokes, trivia, or random facts)
    • Other (Specify)
  3. (4)
    In which ways have you used ChatGPT most frequently? Please order from 1 = most frequent, 2 = next most frequent, and so on.
    • Academic research assistance (e.g., generating citations, summarising articles)
    • Generating text (e.g., drafting e-mails, writing essays or reports)
    • Answering specific questions (e.g., factual, technical, or theoretical questions)
    • Creative writing (e.g., stories, poetry, or brainstorming ideas)
    • Language learning or translation
    • Personal advice or emotional support (e.g., relationship or career advice)
    • Professional or technical problem-solving (e.g., programming, troubleshooting)
    • Entertainment or casual conversation (e.g., jokes, trivia, or random facts)
    • Other (Specify)
Feedback Submission: Thumbs
  1. (5)
    Have you submitted feedback on ChatGPT’s responses using the up and down thumbs?
    • Yes
    • No
  2. (6)
    (If Yes to previous) How often have you submitted feedback using the up and down thumbs?
    • Every time I have used ChatGPT
    • Most times I have used ChatGPT
    • About half the time I have used ChatGPT
    • Some times I have used ChatGPT
    • Rarely
Feedback Submission: Open-Text
  1. (7)
    Have you submitted open-text feedback on ChatGPT’s responses?
    • Yes
    • No
  2. (8)
    (If Yes to previous) How often have you submitted open-text feedback?
    • Every time I have used ChatGPT
    • Most times I have used ChatGPT
    • About half the time I have used ChatGPT
    • Some times I have used ChatGPT
    • Rarely
Feedback Submission: Regenerate Response
  1. (9)
    Have you submitted feedback after regenerating one of ChatGPT’s responses?
    • Yes
    • No
  2. (10)
    (If Yes to previous) How often have you submitted feedback after regenerating a response?
    • Every time I have regenerated a response
    • Most times I have regenerated a response
    • About half the time I have regenerated a response
    • Some times I have regenerated a response
    • Rarely
Reasons for Not Submitting Feedback
  1. (11)
    (If never submitted any feedback) What is the main reason you have not submitted feedback through the ChatGPT interface? Please select the main reason only. NB: ‘Performance’ refers to the helpfulness, truthfulness, or harmlessness of responses from ChatGPT to a prompt.
    • Unaware of feedback option
    • Satisfied with the performance and have not felt the need to submit feedback
    • Unsatisfied with the performance, but do not believe feedback will lead to improvements
    • Unsure how to provide feedback
    • Not enough time or interest in submitting feedback
    • Other (Specify)
Motivations
  1. (12)
    (If submitted open-text feedback) Think back to the most recent time you submitted feedback through the interface. What motivated you to submit feedback about ChatGPT? Please select all that apply.
    • To report a technical issue or bug
    • To suggest new features or improvements
    • To express social or ethical concerns about ChatGPT
    • To comment on use-case specific feedback
    • To contribute to AI development
    • Other (Specify)
  2. (13)
    (If selected multiple motivations previously) What was your main motivation for submitting feedback? Please place a 1 in the box next to your main motivation.
    • To report a technical issue or bug
    • To suggest new features or improvements
    • To express social or ethical concerns about ChatGPT
    • To comment on use-case specific feedback
    • To contribute to AI development
    • Other (Specify)
  3. (14)
    (If submitted open-text feedback) What kind of feedback do you think is most useful for improving ChatGPT’s performance or impact? e.g., reporting technical issues, suggesting new features, expressing social or ethical concerns, and so on. “Performance” refers to the helpfulness, truthfulness, or harmlessness of responses from ChatGPT to a prompt. “Impact” refers to changes in society and culture from ChatGPT. [Open Text]
  4. (15)
    (If submitted open-text feedback) Have you received any follow-up communication or updates regarding your feedback?
    • Yes
    • No
  5. (16)
    (If Yes to previous) How satisfied are you with the response or resolution?
    • Extremely satisfied
    • Somewhat satisfied
    • Neither satisfied nor dissatisfied
    • Somewhat dissatisfied
    • Extremely dissatisfied
  6. (17)
    (If submitted open-text feedback) How much do you agree with this statement: “My reasons for submitting feedback about ChatGPT through the interface have been addressed by OpenAI.”
    • Strongly agree
    • Somewhat agree
    • Neither agree nor disagree
    • Somewhat disagree
    • Strongly disagree
Reflections
  1. (18)
    (If submitted any feedback) Was submitting feedback through the ChatGPT interface easy or difficult?
    • Extremely easy
    • Somewhat easy
    • Neither easy nor difficult
    • Somewhat difficult
    • Extremely difficult
  2. (19)
    How likely are you to submit feedback through the ChatGPT interface in the future?
    • Extremely likely
    • Somewhat likely
    • Neither likely nor unlikely
    • Somewhat unlikely
    • Extremely unlikely
  3. (20)
    (Conditional logic based on Q19) Why are you likely to submit feedback through the ChatGPT interface in the future? [Open Text]
  4. (21)
    (Conditional logic based on Q19) Why are you not likely to submit feedback through the ChatGPT interface in the future? [Open Text]
Feedback via Other Channels
  1. (22)
    Have you provided feedback on ChatGPT’s performance or impact through other channels or platforms? e.g., social media platforms, online discussion forums, contacting OpenAI, and so on. “Performance” refers to the helpfulness, truthfulness, or harmlessness of responses from ChatGPT to a prompt. “Impact” refers to changes in society and culture from ChatGPT.
    • Yes
    • No
  2. (23)
    (If Yes to previous) Which other channels or platforms have you used to provide feedback about ChatGPT’s performance or impact? Please select all that apply.
    • X (Please list any specific # or communities)
    • Mastodon (Please list any specific # or servers)
    • Linkedin (Please list any groups)
    • Facebook (Please list any groups)
    • Reddit forums (Please specify which subreddits e.g., r/ChatGPT)
    • Discord servers (Please specify which servers)
    • Other online discussion forums (Please specify e.g., Github Community, OpenAI Community)
    • Contacting OpenAI (e.g., via e-mail)
    • Other (Specify)
  3. (24)
    (If Yes to using other channels) How likely are you to use those channels or platforms to provide feedback on ChatGPT in the future?
    • Extremely likely
    • Somewhat likely
    • Neither likely nor unlikely
    • Somewhat unlikely
    • Extremely unlikely
  4. (25)
    (Conditional logic based on Q24) Why are you likely to provide feedback on ChatGPT through those channels or platforms in the future? [Open Text]
  5. (26)
    (Conditional logic based on Q24) Why are you not likely to provide feedback on ChatGPT through those channels or platforms in the future? [Open Text]
Demographics
  1. (27)
    What is your age?
    • Under 18
    • 18-24
    • 25-34
    • 35-44
    • 45-54
    • 55-64
    • 65+
    • Prefer not to say
  2. (28)
    How would you describe your gender?
    • Man
    • Woman
    • Non-binary/third gender
    • Prefer to self-describe (Specify)
    • Prefer not to say
  3. (29)
    In which country do you currently reside? [Open Text - Likely]
  4. (30)
    What are the main languages you speak daily? Select up to three languages. [Open Text]
  5. (31)
    Did you experience any barriers accessing ChatGPT or submitting feedback about ChatGPT?
    • Computer hardware barriers
    • Internet connectivity barriers
    • Health and medical barriers
    • Language barriers
    • Other (Specify)
    • None of the above
  6. (32)
    How often do you engage with or read about AI-related topics? e.g., news articles, research papers, podcasts, and so on.
    • Daily
    • Weekly
    • Monthly
    • Rarely
    • Never
  7. (33)
    Before using ChatGPT, have you used any AI, machine learning, or data science tools, beyond basic interaction or exploration? Please select “Yes” if you have used AI, machine learning, or data science tools for a project, work, or in-depth exploration.
    • Yes
    • No
  8. (34)
    Have you completed any formal education, training, or coursework related to AI, machine learning, or data science?
    • Yes
    • No
  9. (35)
    Are you currently working in, or have you ever worked in, a field related to AI, machine learning, or data science?
    • Yes
    • No
  10. (36)
    How much do you agree with the following statement: “I feel confident in my ability to evaluate and provide feedback on the performance and/or impact of generative A systems, such as ChatGPT”?
    • Strongly agree
    • Somewhat agree
    • Neither agree nor disagree
    • Somewhat disagree
    • Strongly disagree
Appendix B Survey Results
Appendix B compiles the survey results from this study.
@@ Appendix B Survey Results
(Response rate: 526/526, Multiple selections allowed).
@@
Every time I have used ChatGPT 23 5.8%
Most times I have used ChatGPT 92 23.0%
About half the time I have used ChatGPT 98 24.5%
@@
Every time I have used ChatGPT 3 2.0%
Most times I have used ChatGPT 20 13.2%
About half the time I have used ChatGPT 17 11.2%
@@
Every time I have regenerated a response 25 7.6%
@@
Satisfied with the performance and have not felt need to submit feedback 17 21.8%
@@
Contacting OpenAI (e.g., via e-mail) 10 9.7%
@@
(Response rate 103/103, Multiple selections allowed).
@@
(Response rate: 509/526, up to 3 selections allowed).
Notes: “Other African languages” includes Setswana, Xhosa, Yoruba, Swahili, and Hausa; “Other European languages” includes Finnish, Hungarian, Latvian, Russian, Norwegian, Danish, Swedish, Catalan, Ukrainian, and Lithuanian; “Other Asian languages” includes Hindi, Hebrew, Turkish, Arabic, Mongolian, Korean, Japanese, Persian, Bengali, Tagalog, Thai, and Sinhalese; “Other languages” includes Portuguese (Brazilian), Slovenian, Croatian, and unspecified others.
@@
(Response rate: 524/526).
Appendix C Statistical Relationships and Test Results
Appendix C presents the statistical analysis of relationships between key variables for this study.
Table 38. Chi-squared Test Results
@@ Appendix C Statistical Relationships and Test Results
Thumbs-Up/down and After Regenerating 50.419 ¡.001 ¡.001* 1 .319 Mod. 497
Education/Training in AI and Confidence 12.773 .002 .023* 2 .160 Weak 497
Working in AI and Barriers 9.821 .002 .023* 1 .143 Weak 477
Region and Open-text feedback 13.990 .003 .023* 3 .192 Mod. 378
Prior AI use and Open-text feedback 8.926 .003 .023* 1 .154 Weak 378
Thumbs-Up/down and ChatGPT frequency 13.689 .008 .045* 4 .166 Weak 497
AI topics frequency and Confidence 20.726 .008 .045* 8 .144 Weak 497
Age and Confidence 17.068 .009 .045* 6 .131 Weak 497
Education/Training in AI and Open-text 6.132 .013 .059 1 .127 Weak 378
Gender and After Regenerating 10.427 .015 .061 3 .145 Weak 497
Age and Barriers 9.372 .025 .078 3 .140 Weak 477
Open-text and ChatGPT frequency 10.956 .027 .078 4 .170 Mod. 378
AI topics frequency and Open-text 10.937 .027 .078 4 .170 Mod. 378
Region and After Regenerating 9.151 .027 .078 3 .136 Weak 497
Working in AI and Open-text feedback 4.660 .031 .080 1 .111 Weak 378
Age and After Regenerating 8.678 .034 .080 3 .132 Weak 497
Prior AI use and Barriers 4.512 .034 .080 1 .097 Negl. 477
Education/Training in AI and Barriers 4.010 .045 .101 1 .092 Negl. 477
Gender and Confidence 12.513 .051 .103 6 .112 Weak 497
Prior AI use and Confidence 5.953 .051 .103 2 .109 Weak 497
Working in AI and Confidence 5.632 .060 .114 2 .106 Weak 497
Open-text and After Regenerating 3.424 .064 .117 1 .095 Negl. 378
AI topics frequency and Thumbs-Up/down 7.615 .107 .186 4 .124 Weak 497
Working in AI and After Regenerating 2.375 .123 .205 1 .069 Negl. 497
Regenerating and ChatGPT frequency 5.004 .287 .459 4 .100 Weak 497
Age and Open-text feedback 3.342 .342 .526 3 .094 Weak 378
Gender and Thumbs-Up/down 3.160 .368 .545 3 .080 Weak 497
AI topics frequency and Barriers 3.962 .411 .554 4 .091 Weak 477
AI topics frequency and Regenerating 3.932 .415 .554 4 .089 Weak 497
Prior AI use and Thumbs-Up/down 0.666 .415 .554 1 .037 Negl. 497
Gender and Barriers 2.482 .478 .607 3 .072 Weak 477
Education/Training in AI and Thumbs 0.468 .494 .607 1 .031 Negl. 497
Working in AI and Thumbs-Up/down 0.453 .501 .607 1 .030 Negl. 497
Table 38. Continued
Relationship $\chi^{2}$ $p$ Adj. $p$ df CV Interp. $n$
Gender and Open-text feedback 1.801 .615 .723 3 .069 Weak 378
Age and Thumbs-Up/down 1.713 .634 .725 3 .059 Negl. 497
Prior AI use and After Regenerating 0.140 .708 .773 1 .017 Negl. 497
Region and Barriers 1.361 .715 .773 3 .053 Negl. 477
Education/Training in AI and Regenerating 0.065 .799 .841 1 .011 Negl. 497
Region and Thumbs-Up/down 0.708 .871 .894 3 .038 Negl. 497
Region and Confidence 1.548 .956 .956 6 .039 Negl. 497
@@
  • $\chi^{2}$ test was not performed between Thumbs-Up/down and Open-text feedback. Open-text feedback was only available to users who provided Thumbs-Up/down feedback, making these variables dependent.
  • Table 39 provides a guide for interpreting Cramér’s V values, in accordance with (Cohen, 2013). Thresholds for df $\geq$ 3 are approximate guidelines.